A plane can fly 480 miles in the same time as it takes a car to go 160 miles. If the car travels 80 mph slower than the plane, find the speed of the plane.

Answer by rfer(16322)

x-80/160=x/480
160x=480x-38400
-320x=-38400
x=120 mph
480/120=160/40
4 hrs= 4 hrs